OpenStudy (anonymous):

Help needed :/ This one is a little different... Find the solutions of the following quadratic equation by completing the square: x^2-x=20* The solutions are: a. -5 or -4 b. -5 or 4 c. 5 or 4 d. 5 or -4

OpenStudy (anonymous):

\[\Large x^2-x=20^* \quad ??? \quad \text{what does * mean??}\]

OpenStudy (anonymous):

*= closed circle= multiply

OpenStudy (anonymous):

\[x ^{2}-5x+4x-20 = 0\] x(x-5) +4(x-5)=0 (x+4)(x-5=0 x=-4,x=5

OpenStudy (anonymous):

does it mean 20 to the power 0 ?? like \[\LARGE 20^0\]

OpenStudy (anonymous):

dont burden him, he hates math, :)

OpenStudy (anonymous):

@myko where does that equation came from ??

OpenStudy (anonymous):

LOL I think myko got it :) Thanks for trying Kreshnik.

OpenStudy (anonymous):

from my keyboard

OpenStudy (anonymous):

Haha :D

OpenStudy (anonymous):

@myko I know it came from your keyboard but if there's \[20^0\] then it's equal to 1 !& @mathsux4real ... I was just trying to know exactly what the given are :( , I wouldn't bother my self asking you if I didn't know the answer would I ?? :)

OpenStudy (anonymous):

that's just a typing mistake. He had been asking questions about quadratic equations all day long...

OpenStudy (anonymous):

but did you read what we write? "*= closed circle= multiply" @myko

OpenStudy (anonymous):

sry @Kreshnik for "from my keyboard" just having a bit of fun

OpenStudy (anonymous):

ok ok .. anyway , sorry for dragging the task down , next time I'll do better :) O_O

OpenStudy (anonymous):

if that would be one, there is no solution

OpenStudy (anonymous):

at least in reals

OpenStudy (anonymous):

@myko Therefore I wanted to be sure, what number it is :(

OpenStudy (anonymous):

kk, sry

OpenStudy (anonymous):

SORRY FOR THE CONFUSION GUYS... :/ Thanks for helping though :)